OpenStudy (anonymous):

Simplify square root of negative 48

OpenStudy (marissalovescats):

Hi and welcome to OpenStudy! As you know, 48 is not a perfect square, so we want to find the largest perfect square that divides evenly into 48. Perfect squares: 1,4,9,16,25,etc So which one is the largest that evenly goes into 48?

OpenStudy (marissalovescats):

If you answer 16, you're correct. 16*3=48 so now under the radical we have this: \[\sqrt{16*3}\] But you ask how do we make i negative? If you recall, i^2=-1 so we do 16*i^2 would be teh same as 16*-1: \[\sqrt{16*i^2*3}\] And we know the square root of 16 is 4, and the square root of i^2 is i. So now we have out final answer: \[4i \sqrt{3}\]

OpenStudy (acxbox22):

\(\Huge\color{green}{\bigstar}\Huge\color{green}{\bigstar}\huge\color{blue}{WELCOME~TO~OPEN~STUDY}\Huge\color{orange}{\bigstar}\Huge\color{orange}{\bigstar}\)

OpenStudy (marissalovescats):

That's really cool lol

OpenStudy (acxbox22):

are you looking for real roots or imaginary roots?

OpenStudy (acxbox22):

if real then there are no real solutions to the sqrt of -48

OpenStudy (acxbox22):

thx @marissalovescats

OpenStudy (marissalovescats):

Yes no real ones, but imaginary ones as I put